I am needing some help figuring out the blanks. I have been working on this problem for a couple of hours and was hoping someone can walk me through it. Thank YouExercise 4-26 (Algorithmic) (LO. 4) Determine the taxable amount of social security benefits for the following situations. If
Answer #1

For married filing jointly, the social security benefits are taxable to 50% for the combined income between $32,000 and $44,000.

For married filing jointly, the social security benefits are taxable to 85% for the combined income more than $44,000.

Part A

Taxable amount of social security benefits is lower of the following two:

Taxable amount of social security benefits = 14560*50% = $7280

Taxable amount of social security benefits = 41600+(14560*50%)-32000 = $16880

Therefore,

Taxable amount of social security benefits = $7280

Part B

Taxable amount of social security benefits is lower of the following two:

Taxable amount of social security benefits = 18040*50% = $9020

Taxable amount of social security benefits = 16400+(18400*50%)-32000 = -6440 = $0

Therefore,

Taxable amount of social security benefits = $0

Part C

Taxable amount of social security benefits is lower of the following two:

Taxable amount of social security benefits = 15000*85% = $12750

Taxable amount of social security benefits = 100000+(15000*85%)-44000 = $6750

Therefore,

Taxable amount of social security benefits = $12750